sir, i have asked your help for calculating the ristance of a conical conductor. but still ihave a confusion. the area of cross section will be not same so what will be the value of A?kindly help me again. i would be obligue to you.

Grade:

1 Answers

vikas askiitian expert
13 years ago

let the radius of base is R & height of cone is H ........(given)

now place the base in XY plane with center as origin so that the top most point of cone lies on Z axis....

semi verticle angle of cone = tan@ = R/H  ..............1

take a small elemental disk whose thickness is dz at a distance Z from origin....

radius of disc is r ...

from here tan@ = r /H-Z       ............2

equating 2 & 3

r/H-Z = R/H   ..........3

now find the resistance for this elementl disc & integrate it using proper limits...